OpenStudy (stamp):

[CALCULUS] Determine the area of the region bounded by the curves f and g for -1.5 ≤ x ≤ 0.5

OpenStudy (stamp):

\[f=cos(x^2)\]\[g=e^x\] -1.5 ≤ x ≤ 0.5

OpenStudy (stamp):

I just need somebody to crunch the number for me. I have my answer and want to be sure I explained it right to somebody else.

OpenStudy (stamp):

Intercepts -1.11 and 0

OpenStudy (legomyego180):

\[\int\limits_{-1.5}^{0}\cos(x^2)-e^xdx+\int\limits_{0}^{.5}e^x-\cos(x^2)dx\]

OpenStudy (legomyego180):

Thats how I would do it, you might be able to do it with one integral instead of two but I wasnt 100% sure.

OpenStudy (eliesaab):

You need to do numerical Integration to compute the integral involving \( \cos^2(x)\)

OpenStudy (eliesaab):

You have to compute three integrals \[\int_{-1.5}^{-1.11} \left(e^x-\cos \left(x^2\right)\right) \, dx+ \int_{-1.1}^0 \left(\cos \left(x^2\right)-e^x\right) \, dx+ \int_0^{0.5} \left(e^x-\cos \left(x^2\right)\right) \, dx \]

OpenStudy (eliesaab):

All the three needs to be done with Numerical Integrtion

OpenStudy (eliesaab):

\[ \int_{-1.5}^{-1.11} \left(e^x-\cos \left(x^2\right)\right) \, dx=0.160175 \]

OpenStudy (eliesaab):

\[ \int_{-1.1}^0 \left(\cos \left(x^2\right)-e^x\right) \, dx=0.282375 \]

OpenStudy (eliesaab):

Do the third integral yourself

OpenStudy (stamp):

Like I said, I already did this. I was wondering what yall got for the final. 0.59

OpenStudy (stamp):

I helped somebody with this problem earlier. I want to verify the answer we got A = 0.59 is correct, if not I will send them the updates.

OpenStudy (eliesaab):

0.594387

OpenStudy (mathmale):

Remember that you're approx. an area. Thus, giving an answer to 6 decimal place accuracy is stretching things quite a bit. What do you think would be sufficient accuracy for an area approximation?